OpenStudy (anonymous):

discriminant: COMPUTE THE DISCRIMINANT: (d =b^2 -4ac) for the equation: (-1/6x^2) +3x -9 = 0

OpenStudy (anonymous):

9-4(-1/6)(-9)=9-36/6=9-6=3

OpenStudy (anonymous):

how do you solve can i get the steps :)

OpenStudy (anonymous):

eitaK: Check out the following: http://answers.yahoo.com/question/index?qid=20080304153003AAUaVrd

OpenStudy (radar):

Hey eitaK, I would multiply all terms by -6 to get: \[x ^{2}-18x+54=0\] then the discriminant would be \[-18^{2}-(4)(1)(54)=324-216=108\]

OpenStudy (anonymous):

so technically you are getting rid of the fraction in the first step right?

OpenStudy (anonymous):

and then how did you do the next step?

OpenStudy (radar):

Yesm I believe twhitelsu is also correct. I am sort of confused here.

OpenStudy (radar):

The next step was to square b (-18*-18) then subtract the 4ac or subtract (4*1*54) The results did not equal the same as twhitelsu, and there is nothing wrong with the way it was done to get 3

OpenStudy (anonymous):

ok but what happened to the x's?

OpenStudy (radar):

I would stick the a, b, c, values of the original equation

OpenStudy (anonymous):

ok so that is what i didn't and i go t -18^2- 4(1)(54)

OpenStudy (anonymous):

sorry d= -18 not t.

OpenStudy (radar):

The discriminant only deals with the coefficients of the variable not the variable itself. So there are no x in the discriminant.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so then i went further and got-18^2 = -324 and then -4(1)(54) = -216

OpenStudy (anonymous):

so then i woudl have D= -324 -216 correct?

OpenStudy (radar):

Yes, but it doesn't come out as 3, as you would get if you had used the original equation for the values of a, b, c. I thought multiplying thru would simplify it, but it came out with a different discriminant. So apparently multipllying thru by-6 was not the right thing to do. I will think some more about this. I think you are on to the procedure to obtain discriminants.

OpenStudy (anonymous):

i just used the equation that i was given. that i put abvoe :/

OpenStudy (radar):

I just noticed that if I divides the results of 108 by 6 squared (36) it would be 3!!.

OpenStudy (anonymous):

wait but where did you get 108

OpenStudy (anonymous):

you've lost me,

OpenStudy (radar):

The discriminant is that part of the quadratic formula that is within the radical and it is used to the following way: If it is greater than 0, their are two real solutions. If it equals zero there is one real solution (repeated). If it is less than 0, there are no real solutions.

OpenStudy (radar):

The -18 squared results in positive 324 (not the -324) when then combined with the -216 you get the 108 which is positive and means there are 2 real solutions.

OpenStudy (anonymous):

ok... because it it positive?

OpenStudy (anonymous):

and greater than 0 correct?

OpenStudy (radar):

Yes

OpenStudy (anonymous):

ok so then how do i get three as the discriminant?

OpenStudy (radar):

Look at the work at the top of this page provided by twhitelsu.

OpenStudy (radar):

She worked the probllem using the original equation working with the fractional value for a.

OpenStudy (anonymous):

if i am going of yours, which makes a little more sense to me... how did you think and know to do that?

OpenStudy (radar):

Did you follow her steps, see where b (3) was squared to become a 9

OpenStudy (anonymous):

no i don't see where she got the 4 from.

OpenStudy (radar):

I was trying to get rid of the 1/6 (I don't like fractions) lol the result were still positive but was not equal

OpenStudy (anonymous):

i mean looking at hers. i see she has 4

OpenStudy (radar):

the 4 was given in the discriminant equation: b^2-4ac, it is part of the 4ac. get it

OpenStudy (radar):

you can see where she uses the a and c values being that a = -1/6 and c=-9

OpenStudy (anonymous):

oh yes haha thanks.

OpenStudy (anonymous):

but isn't the b supposed to be squared ? so 9^2

OpenStudy (radar):

Look at your equation (-1/6)x^2+3x-9=0 the abc values for your equation are: a=-1/6, b=3, c=-9 So b=3 and b squared if 9

OpenStudy (anonymous):

gotcha! wow haha

OpenStudy (anonymous):

THANKS FOR SPENDING ALL THAT TIME ON THIS PROB!

OpenStudy (radar):

Here is form for a quadratic equation: \[ax ^{2}+bx+c=0\] Note that the abc represents numbers with a and b being the coefficients of variables and c is a constant.

OpenStudy (anonymous):

wait whats that for?

OpenStudy (anonymous):

oh i see.